# The Gaussian-Multinoulli Restricted Boltzmann Machine A Potts Model Extension of the GRBM

### v1

@@ 4 Hetero-associative Memory
Hetero-associative memory refers to a system’s capability to learn paired associations between distinct patterns, allowing the retrieval of a target pattern (e.g., a response word) when presented with a corresponding stimulus (e.g., a cue word) [11, 19]. This concept, rooted in cognitive modeling and neural computation, was further explored by Hinton in 1981 [2] and later employed in language-related tasks using Gaussian-Bernoulli Restricted Boltzmann Machines (GB-RBMs) [27]. However, the binary hidden units of GB-RBMs impose a hard ceiling on representational capacity.
@@ 4.1 Experimental Setup
4.1.1 Dataset
We replicated the experimental setup in Tsutsui and Hagiwara, constructing a word-pair dataset representing conceptual relationships (e.g., "apple is-a fruit") [27]. We randomly selected pairs from WordNet [28], excluding compound or incomplete entries and sampled 500-3,000 word pairs to create smaller datasets for training and testing scalability.
Each word pair is treated as a directional association task and encoded as a concatenated embedding vector [17], mirroring the hetero-associative memory objective of recalling a target concept from a given stimulus.
4.1.2 Preprocessing
We trained a 200-dimensional Continuous Bag of Words (CBOW) Word2Vec model [18] on the word-pair dataset (100 iterations, window size 5, no frequency cutoff). We normalized each vector dimension to zero mean and unit variance to mitigate small-magnitude issues. For each stimulus-response pair, embeddings were concatenated to form a 400-dimensional input vector for the RBM’s visible layer.
4.1.3 Computational Setup
@@
4.1.4 Training
Our GM-RBM learned to associate stimulus–response word embeddings. Datasets comprised semantically related word pairs (e.g., doctor–nurse, sun–light). Following a similar procedure to Tsutsui and Hagiwara [27], we trained a CBOW Word2Vec model [23] on these two-word sentences to capture in-domain semantics, producing 200-dimensional embeddings (100 iterations, window size 5, no frequency cutoff).
@@ 4.2 Results
To isolate the effect of the Potts hidden units, we did two key tests. The first was a parameter-matched comparison where the total number of weights was kept equal while $q$ was increased. This ensured that any observed gains in recall accuracy arise purely from the richer Potts representation. The second was a similar sweep where $q$ was held constant while the number of hidden units themselves was increased. Both sweeps were done across varying dataset sizes, measuring the accuracy of the recall in pairs embedded in Word2Vec [18].
@@
It is also important to note $q=2$ for the GM-RBM is a case of the Potts nodes where weights do not have to be constrained i.e. negatives of one another. The original GB-RBM outperforms the GM-RBM for $q=2$, possibly due to the lack of this constraint. Despite the GM-RBMs being trained using the simpler Gibbs update, the $q=4,6,8,10$ GM-RBM models all vastly outperform the GB-RBM, which used the more expensive Gibbs-Langevin update [15].
@@ 4.2.1 Parameter matched q sweep
To keep the model capacity fixed across different Potts-state configurations, we solve for the number of hidden units $n_{h}$ by dividing the total budget of weight parameters $n_{w}$ by the size of each hidden Potts spin $q$. Since each hidden unit with $q$ Potts states contributes $q$ weight vectors of length equal to the number of visible units $n_{v}$, then number of hidden units is given by
@@
We chose $\mathrm{n_{w}}=800{,}000$, ensuring that the binary case $q=2$ matched the optimal hidden-unit count of 1,000 reported by Tsutsui and Hagiwara [27]for the GB-RBM. This ensures that, as we increase the number of Potts states $k$, we reduce the number of hidden units proportionally so that every model has the same total number of parameters.
@@ 4.2.2 Hidden nodes sweep
Figure 2: Retrieval accuracy on a semantic hetero-associative memory task as a function of hidden-layer size for different model variants and dataset sizes. (a) GB-RBM (Gibbs–Langevin Update), (b) GM-RBM with $q=2$ (Gibbs Update), and (c) GM-RBM with $q=4$ (Gibbs Update). Each curve corresponds to a different number of training word-pair examples ($N$).
@@ 5 Auto-associative memory
As a "proof-of-concept" to demonstrate the generative capability of our GM-RBM, we performed a replication of the generative experiments from the original GB-RBM paper [15]. On MNIST, we trained for 500 epochs, and on CelebA for 100 epochs. Starting from i.i.d. Gaussian noise in the visible layer and running 1,000 steps of Gibbs sampling, we obtained samples (see Figure 3) that, by visual inspection, showcasing the GM-RBM’s ability to still produce high-quality generative outputs.
@@ 5.1 Experimental setup
5.1.1 Datasets
We target two key datasets to showcase the GM-RBM’s generative abilities: MNIST (28×28 grayscale handwritten digits) [13] and CelebA (Center-cropped and resized RGB facial images at 64×64 resolution)  [16].
5.1.2 Preprocessing
All image datasets are normalized to zero mean and unit variance per channel [15]. CelebA images are center-cropped to 140×140 before downsampling to 64×64 [16]. The GMM datasets are constructed using predefined priors and covariance structures, allowing for exact density visualization and sampling [15]. For each dataset, we generate training and evaluation splits using standard protocols.
5.1.3 Computational resources
@@ 5.2 Results
(b) Generation of CelebA images from random noise
Figure 3: Sampled results from GM-RBM of MNIST and CelebA datatsets
We observe that the $q=4$ GM-RBM begins to generate visually identifiable face/digit samples with an order of magnitude lower number of epochs compared to the GB-RBM trained with Gibbs–Langevin sampling (See Table 2. Although a fully controlled head-to-head comparison under identical hyperparameter budgets (see Table 2) requires further investigation (See Section 6.1, it is important to emphasize two key distinctions. Our GM-RBM relies solely on conventional Gibbs sampling, whereas Hinton’s GB-RBM employs the more computationally expensive Gibbs–Langevin procedure [15]. This ability is most likely due to what the rapid mixing of Potts models compared to binary Boltzmann machines [22, 31, 32]. In addition, we maintain the same total number of parameters while increasing the number of Potts states.
@@
Hyperparameter MNIST CelebA
GB-RBM  [15] GM-RBM GB-RBM [15] GM-RBM
@@ 6 Limitations and future work
While the Gaussian-Multinoulli Restricted Boltzmann Machine (GM-RBM) demonstrates improved expressiveness and recall performance over the binary GB-RBM, key limitations and future work remain.
@@ 6.1 Limitations of the current model
1. Sampling constraints: In the current implementation, only Gibbs sampling is supported for the Potts-based GM-RBM. While this allows for tractable inference over discrete states, the absence of Langevin or hybrid Gibbs-Langevin samplers limits potential gains in sample quality and mixing speed. As $q$ increases, the discrete state space grows, and relying solely on categorical sampling may lead to slower convergence or poor exploration of the energy landscape in more complex settings.
2. Evaluation domain: The current evaluation focuses on word pair associations using Word2Vec embeddings. The generalization of GM-RBMs to other domains, such as vision, audio, or time series data, remains an open question and will require further validation. These limitations present valuable directions for future research, including scaling GM-RBMs to deeper architectures, optimizing discrete sampling methods, and exploring broader applications beyond associative memory tasks.
3. Epochs and training: Currently, the GM-RBM is trained for a limited number of epochs (500 for MNIST, 100 for CelebA) to achieve initial convergence. While this rapid training demonstrates efficiency, it may not capture long-term stability or the potential benefits of extended training schedules. Optimal epoch budgets, early-stopping criteria, and the impact of prolonged training on sample diversity and model robustness are areas for further exploration.
@@ 6.2.1 Potts units in energy transformers
Energy Transformers (ETs) minimize a global energy via recurrent updates and have shown success in tasks such as image inpainting [9]. Replacing their binary hidden units with $q$-state Potts variables raises the latent capacity from $2^{H}$ to $q^{H}$, yielding:
$E_{\mathrm{mem}}(h)=-\sum_{\mu=1}^{M}\sum_{i=1}^{H}\delta(h_{i},\xi_{i}^{\mu})$ (6)
where ${\xi^{\mu}}$ are stored patterns. Larger $q$ reduces overlap between attractors, improving convergence selectivity. Embedding Gaussian–Potts RBMs into ETs may boost robustness and facilitate neuromorphic implementations; an empirical study of these hybrids is ongoing.
6.2.2 Extending DBMs to continuous data
Deep Boltzmann Machines (DBMs) stack Bernoulli–Bernoulli RBMs to learn hierarchical features via layer-wise Gibbs sampling and contrastive divergence [3, 24]. Their reliance on binary visibles limits real-valued inputs. As a front end, we propose a Gaussian–Potts energy:
$E(v,h)=\sum_{i=1}^{D}\frac{(v_{i}-a_{i})^{2}}{2\sigma_{i}^{2}}-\sum_{i,j}W_{ij}(v_{i},h_{j})-\sum_{j}b_{h_{j}},$ (7)
with $h\_j\in{1,\dots,q}$. This encoder maps $\mathbb{R}^{D}$ into a $q^{H}$-state latent space, preserving DBM training while improving stability and expressiveness.
6.3 Bernoulli-Variable models and Potts benefits
Many foundational generative architectures rely on binary or Bernoulli units, including classical models like Bernoulli RBMs [6], Gaussian–Bernoulli RBMs [15], Deep Belief Networks [24], Deep Boltzmann Machines [3, 24], and Hopfield Networks. More recent advancements leveraging binary sampling mechanisms include Variational Autoencoders (VAEs) [5] with Bernoulli latents, Gumbel–Softmax VAEs [21] that provide differentiable relaxations for discrete sampling, Autoregressive Models such as PixelCNN [30] and PixelRNN [29] for sequential binary pixel generation, Binarized Neural Networks (BNNs) that constrain weights and activations to binary values, Latent Bernoulli Autoencoders [5] designed explicitly for compact binary representations, Spiking Neural Networks (SNNs) employing discrete spike events for communication, and Discrete Denoising Diffusion Models (D3PMs) [8] that extend diffusion processes to discrete data through multinomial transition kernels. Upgrading these binary units to Potts variables further increases representational granularity and sharpens attractor basins, significantly reducing interference among stored patterns and enabling richer discrete structures, ultimately enhancing the capability and expressiveness of associative and generative models.
6.4 Hardware-Efficient Implementation
Many of these models can be deployed with minimal overhead by leveraging sparse or event-driven representations (e.g., SPAD arrays [31]) and one-hot encoding of discrete states. Binary and Potts variables map naturally to bitwise logic and lookup tables, enabling FPGA, ASIC, and neuromorphic implementations with low memory footprint and energy consumption.

### v2

@@ 4 Hetero-associative Memory
Hetero-associative memory refers to a system’s capability to learn paired associations between distinct patterns, allowing the retrieval of a target pattern (e.g., a response word) when presented with a corresponding stimulus (e.g., a cue word) (Kosko, 1988; Morales and Pineda, 2024). This concept, rooted in cognitive modeling and neural computation, was further explored by Hinton in 1981 Anderson and Hinton (1981) and later employed in language-related tasks using Gaussian-Bernoulli Restricted Boltzmann Machines (GB-RBMs) (Tsutsui and Hagiwara, 2019). However, the binary hidden units of GB-RBMs impose a hard ceiling on representational capacity.
@@ 4.1 Experimental Setup
We replicated the experimental setup in Tsutsui and Hagiwara, constructing a word-pair dataset representing conceptual relationships (e.g., "apple is-a fruit") (Tsutsui and Hagiwara, 2019). We randomly selected pairs from WordNet (Princeton University (2010)), excluding compound or incomplete entries and sampled 500-3,000 word pairs to create smaller datasets for training and testing scalability.
Each word pair is treated as a directional association task and encoded as a concatenated embedding vector (Mikolov et al., 2013), mirroring the hetero-associative memory objective of recalling a target concept from a given stimulus.
We trained a 200-dimensional Continuous Bag of Words (CBOW) Word2Vec model (Mikolov et al., 2013) on the word-pair dataset (100 iterations, window size 5, no frequency cutoff). We normalized each vector dimension to zero mean and unit variance to mitigate small-magnitude issues. For each stimulus-response pair, embeddings were concatenated to form a 400-dimensional input vector for the RBM’s visible layer.
@@
4.2 Training
Our GM-RBM learned to associate stimulus–response word embeddings. Datasets comprised semantically related word pairs (e.g., doctor–nurse, sun–light). Following a similar procedure to Tsutsui and Hagiwara Tsutsui and Hagiwara (2019), we trained a CBOW Word2Vec model (Řehůřek and Sojka, 2010) on these two-word sentences to capture in-domain semantics, producing 200-dimensional embeddings (100 iterations, window size 5, no frequency cutoff).
@@ 4.3 Results
To isolate the effect of the Potts hidden units, we did two key tests. The first was a parameter-matched comparison where the total number of weights was kept equal while $q$ was increased. This ensured that any observed gains in recall accuracy arise purely from the richer Potts representation. The second was a similar sweep where $q$ was held constant while the number of hidden units themselves was increased. Both sweeps were done across varying dataset sizes, measuring the accuracy of the recall in pairs embedded in Word2Vec (Mikolov et al., 2013).
@@
It is also important to note $q=2$ for the GM-RBM is a case of the Potts nodes where weights do not have to be constrained i.e. negatives of one another. The original GB-RBM outperforms the GM-RBM for $q=2$, possibly due to the lack of this constraint. Despite the GM-RBMs being trained using the simpler Gibbs update, the $q=4,6,8,10$ GM-RBM models all vastly outperform the GB-RBM, which used the more expensive Gibbs-Langevin update (Liao et al., 2022).
@@ 4.3.1 Parameter matched q sweep
To keep the model capacity fixed across different Potts‐state configurations, we solve for the number of hidden units $n_{h}$ by dividing the total budget of weight parameters $n_{w}$ by the size of each hidden Potts spin $q$. Since each hidden unit with $q$ Potts states contributes $q$ weight vectors of length equal to the number of visible units $n_{v}$, then number of hidden units is given by
@@
We chose $\mathrm{n_{w}}=800{,}000$, ensuring that the binary case $q=2$ matched the optimal hidden‐unit count of 1,000 reported by Tsutsui and Hagiwara Tsutsui and Hagiwara (2019)for the GB-RBM. This ensures that, as we increase the number of Potts states $k$, we reduce the number of hidden units proportionally so that every model has the same total number of parameters.
@@ 4.3.2 Hidden nodes sweep
Figure 2: Retrieval accuracy on a semantic hetero‐associative memory task as a function of hidden‐layer size for different model variants and dataset sizes. (a) GB-RBM (Gibbs–Langevin Update), (b) GM-RBM with $q=2$ (Gibbs Update), and (c) GM-RBM with $q=4$ (Gibbs Update). Each curve corresponds to a different number of training word‐pair examples (NN).
@@ 5 Auto-associative memory
As a "proof-of-concept" to demonstrate the generative capability of our GM-RBM, we performed a replication of the generative experiments from the original GB-RBM paper Liao et al. (2022). On MNIST, we trained for 500 epochs, and on CelebA for 100 epochs. Starting from i.i.d. Gaussian noise in the visible layer and running 1,000 steps of Gibbs sampling, we obtained samples (see Figure 3) that, by visual inspection, showcasing the GM-RBM’s ability to still produce high-quality generative outputs.
@@ 5.1 Experimental setup
We target two key datasets to showcase the GM-RBM’s generative abilities: MNIST (28×28 grayscale handwritten digits) (LeCun et al., 1998) and CelebA (Center-cropped and resized RGB facial images at 64×64 resolution)  (Liu et al., 2015).
All image datasets are normalized to zero mean and unit variance per channel (Liao et al., 2022). CelebA images are center-cropped to 140×140 before downsampling to 64×64 (Liu et al., 2015). The GMM datasets are constructed using predefined priors and covariance structures, allowing for exact density visualization and sampling (Liao et al., 2022). For each dataset, we generate training and evaluation splits using standard protocols.
@@ 5.2 Results
(b) Generation of CelebA images from random noise with $q=4$
Figure 3: Sampled results from GM-RBM of MNIST and CelebA datasets.
We observe that the $q=4$ GM-RBM begins to generate visually identifiable face/digit samples (See Figure 3(a) and 3(b)) with an order of magnitude lower number of epochs compared to the GB-RBM trained with Gibbs–Langevin sampling (See Table 2. Although a fully controlled head-to-head comparison under identical hyperparameter budgets (see Table 2) requires further investigation (See Section 6.1), it is important to emphasize two key distinctions. Our GM-RBM relies solely on conventional Gibbs sampling, whereas Hinton’s GB-RBM employs the more computationally expensive Gibbs–Langevin procedure (Liao et al., 2022). This ability is most likely due to the rapid mixing of Potts models compared to binary Boltzmann machines (Potts, 1952; Whitehead et al., 2023; Wu, 1982). In addition, we maintain the same total number of parameters while increasing the number of Potts states.
@@
MNIST CelebA
Hyperparameter GB-RBM GM-RBM GB-RBM GM-RBM
@@
5.2.1 Quantitative sample quality (FID)
Table 3: FID (↓) under parameter-matched budgets. Best overall in bold.
Model Potts States FID
GM–RBM $q=2$ 67.08
GM–RBM $q=4$ 56.09
GM–RBM $q=6$ 53.07
GB–RBM N/A 60.06
To complement the qualitative samples, we report Fréchet Inception Distance (Heusel et al., 2018) (FID; ↓ is better) under a capacity-matched protocol ($m^{\prime}\approx m\log_{2}q$), holding negative-phase budgets fixed across models. Results show that increasing $q$ consistently improves GM–RBM sample quality under a pure Gibbs sampler, narrowing the gap to GB–RBM with Gibbs–Langevin despite the latter’s higher per-step compute.
With the Gibbs-only updates, GM–RBM (q=6q{=}6) outperforms GB–RBM by $\approx$7 FID points (53.07 vs 60.06). With a more expensive update, it is possible to lower the FID score. It is important to note that the CelebA dataset does not necessarily have clear relational qualities, meaning that these scores are really showcasing the mixing ability of the different sampling algorithms and models.
@@ 6 Limitations and future work
While the Gaussian–Multinoulli RBM (GM–RBM) demonstrates stronger discrete expressiveness and robust recall under fair matching, important limitations and next steps remain.
@@ 6.1 Limitations of the current model
1. Sampling constraints and rationale. Our GM–RBM uses pure block Gibbs (exact Gaussian visible draw + per-slot softmax posteriors). We intentionally avoid visible-space Langevin during training because it adds step-size hyperparameters, extra updates, and discretization error, increasing compute relative to a single exact Gaussian draw.
2. Evaluation scope and scaling. Most of our positive results are on hetero-associative recall with in-domain Word2Vec embeddings and proof-of-concept image generation. This leaves open: (a) robustness to stronger or off-the-shelf embeddings (e.g., large CBOW/skip-gram trained on broader corpora), (b) transfer to other modalities (audio, trajectories, time series), and (c) deeper stacks (e.g., GM front-ends for DBMs/DBNs). We also observe retrieval degradation as $N$ grows in some settings; diagnosing whether this arises from chain mixing, capacity-matching choices, or dataset effects will require ablations (burn-in length, persistence, $q$ vs. $m$, negative-phase steps).
3. Training budget and stability. Our generative experiments used short schedules (e.g., 500 epochs on MNIST, 100 on CelebA) to highlight practicality under a reduced budget. Longer training with early-stopping and variance monitoring could reveal stability regimes, impacts on sample diversity, and when heavier negative-phase samplers pay off.
@@ 6.2.1 Potts units in energy transformers
Energy Transformers minimize a global energy via recurrent updates. Replacing binary hidden units with $q$-state Potts slots increases latent capacity from $2^{H}$ to $q^{H}$ and reduces attractor overlap:
$E_{\mathrm{mem}}(h)\;=\;-\sum_{\mu=1}^{M}\sum_{i=1}^{H}\delta(h_{i},\xi_{i}^{\mu}).$ (2)
We hypothesize improved selectivity/robustness under higher $q$ and are empirically exploring GM–RBM components as ET memories; careful comparisons will match either capacity or parameter budgets as above. (Schröder et al., 2024; Hoover et al., 2023)
As a front-end to DBMs, a Gaussian–Potts encoder maps $\mathbb{R}^{D}$ into a $q^{H}$ latent space while preserving layer-wise Gibbs training. Future work should examine stability under deeper stacks and quantify benefits over increased Bernoulli width at matched capacity. (Morales and Pineda, 2024; Oh et al., 2022)
Many foundational generative families, RBMs, DBNs/DBMs, Hopfield-style memories, VAEs with Bernoulli latents, Gumbel-Softmax relaxations, autoregressive binarized pixel models, BNNs, SNNs, and discrete diffusion—lean on binary sampling. Swapping binary units for Potts slots increases representational granularity and sharpens attractor basins, which can reduce interference among stored patterns at equal capacity. A broad, capacity- or parameter-matched survey across these families is an open, high-impact direction (Kosko, 1988).
Binary and Potts one-hot codes map naturally to LUTs and bitwise logic. Sparse/event-driven readout and lightweight on-chip softmax enable efficient FPGA/ASIC/neuromorphic realizations; exploring SPAD-assisted annealing or mixed-signal implementations for categorical slots is promising future work (Whitehead et al., 2023).